OpenStudy (anonymous):

i need help again ! can anyone help ? What is the percentile for data value 6 in the following data set? 4 13 8 6 4 4 13 6 4 13 2 13 15 5 9 4 12 8 6 13 40 35 25 62

OpenStudy (anonymous):

one thing... Didn't get the question yet..../// :/ Specific please...

OpenStudy (anonymous):

@harishk idk how to i copied and pasted

OpenStudy (anonymous):

Order the values: 2 4 4 4 4 4 5 6 6 6 8 8 9 12 13 13 13 13 13 15 There are 20 values. 7 values are less than 6.

OpenStudy (anonymous):

Can you tell me what you think the percentile rank is?

OpenStudy (anonymous):

like the definition ?

OpenStudy (anonymous):

no, the number :)

OpenStudy (anonymous):

would it be 12 ?

OpenStudy (anonymous):

no, 6

OpenStudy (anonymous):

so it would be the median ?

OpenStudy (anonymous):

So what would the percentile rank of 6 be?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

i can give you an example with a different number

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Percentile rank of 14 is (9/15)(100) = 60 or 60th percentile

OpenStudy (mrnood):

@pinkbubbles why 9/15? 12/20 is more intuitive (I acknowledge it is the same number...)

OpenStudy (anonymous):

Okay! :) @MrNood

OpenStudy (anonymous):

ugh im sorry i dont get this at all !

OpenStudy (anonymous):

it's fine! :D

OpenStudy (anonymous):

\[\frac{ number of scores below x }{ n }\times100 = percentile rank\] So x is 6 andnumber of scores below 6 is 7

OpenStudy (anonymous):

so 7(6)/n(100)=percentile rank

OpenStudy (anonymous):

\[\frac{ 7 }{ 6 }\times100=? \]

OpenStudy (anonymous):

Correct! :D

OpenStudy (anonymous):

116.67 ?

OpenStudy (anonymous):

but that isnt one of the choices

OpenStudy (anonymous):

Oops sorry i made a mistake so sorry

OpenStudy (anonymous):

no its fine

OpenStudy (anonymous):

\[\frac{ 7 }{ 20 }\times100=\]

OpenStudy (anonymous):

so 35 ?

OpenStudy (anonymous):

Right! =)

OpenStudy (anonymous):

Do you understand the formula now?

OpenStudy (anonymous):

@dani_00

OpenStudy (anonymous):

oh sorry ! yes i do thank you !!

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (anonymous):

Welcome! If you need anymore help i'm here just tag me! :D

OpenStudy (anonymous):

okay thank you !